Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

A hand purse contains 6 nickels, 5 pennies and 4 dimes.

Expert replies
by netigen » Thu Jun 12, 2008 11:12 pm
A hand purse contains 6 nickels, 5 pennies and 4 dimes. What is the probability of picking a coin other than a nickel twice in a row if the first coin picked is not put back?
Join the discussion
Source: — Problem Solving |

netigen wrote:A hand purse contains 6 nickels, 5 pennies and 4 dimes. What is the probability of picking a coin other than a nickel twice in a row if the first coin picked is not put back?
Probability of picking a P or D in first pick = (5+4)/15 = 9/15 = 3/5
Probability of picking a P or D in second pick = 8/14 = 4/7

Combined probability = 3/5 * 4/7 = 12/35
Join the discussion

by netigen » Thu Jun 12, 2008 11:35 pm
You are assuming that the Q means we have to pick any non nickel coin twice.

I read the question as saying that the same non nickle coin has to be picked twice.
Join the discussion

by durgesh79 » Thu Jun 12, 2008 11:47 pm
In that case the answer will be

P : 5/15 * 4/14
D : 4/15 * 3/14

Final probability = (20 + 12) / 15*14 = 16/105

The answer will depend on the options available . If both 12/35 and 16/105 are options, then its a tricky question. :?
Join the discussion

by netigen » Thu Jun 12, 2008 11:51 pm
Question is can this question be interpreted in multiple ways or am I just over reading :)
Join the discussion

by jasonc » Fri Jun 13, 2008 12:47 am
you're overreading :)
that, or its not an OG question :p
I beat the GMAT! 760 (Q49/V44)
Join the discussion

by durgesh79 » Fri Jun 13, 2008 9:41 am
May be I'm overthinking but looks like this one is an actual GMAT problem. Refer below thread, why auther would think that he might be in trouble. :wink:

https://www.beatthegmat.com/probability-t10305.html
Join the discussion

by netigen » Fri Jun 13, 2008 12:30 pm
This is a MGMAT Q
Join the discussion

by Scott@TargetTestPrep » Wed Nov 20, 2019 5:53 pm
netigen wrote:A hand purse contains 6 nickels, 5 pennies and 4 dimes. What is the probability of picking a coin other than a nickel twice in a row if the first coin picked is not put back?
The probability of picking a coin that is not a nickel twice in a row is 9/15 * 8/14 = 3/5 * 4/7 = 12/35.

Scott Woodbury-Stewart
Founder and CEO
[email protected]

Image

See why Target Test Prep is rated 5 out of 5 stars on BEAT the GMAT. Read our reviews

ImageImage
Join the discussion

by [email protected] » Thu Dec 12, 2019 12:22 pm
Hi All,

We're told that a hand purse contains 6 nickels, 5 pennies and 4 dimes. We're asked for the probability of picking two coins OTHER than nickels if the first coin picked is NOT put back. This question is a straight-forward Probability question, so we just have to work step-by-step and do the necessary calculations...

There are 6+5+4 = 15 total coins.

The probability of NOT choosing a nickel on the first try is 9/15 = 3/5

Since we do NOT put that coin back, we have removed one non-nickel from the purse and the probability of NOT choosing a nickel on the second try is 8/14 = 4/7

Thus, the overall probability of pulling two NON-nickels is (3/5)(4/7)...

Final Answer: [spoiler]12/35[/spoiler]

GMAT assassins aren't born, they're made,
Rich
Contact Rich at [email protected]
Image
Join the discussion